OpenStudy (anonymous):

The temperature on a summer afternoon is 76°F. Define a variable and write an expression to find the temperature after it changes. Then evaluate your expression for a decrease of 8 degrees Fahrenheit.

OpenStudy (mathstudent55):

Let the temperature = t After it changes, it becomes t - 8 T = current temperature T = t - 8 T = 76 - 8 T = 68